# On the Existence of Steady-State Solutions to the Equations Governing Fluid Flow in Networks

### 23 Jan 2024 v3

@@ Abstract
The steady-state solution of fluid flow in pipeline infrastructure networks driven by junction/node potentials is a crucial ingredient in various decision support tools for system design and operation. While the non-linear system is known to have a unique solution (when one exists), the absence of a definite result on existence of solutions hobbles the development of computational algorithms, for it is not possible to distinguish between algorithm failure and non-existence of a solution. In this letter we show that a unique solution exists for such non-linear systems if the term solution is interpreted in terms of potentials and flows rather than pressures and flows. The existence result for flow of natural gas in networks also applies to other fluid flow networks such as water distribution networks or networks that transport carbon dioxide in carbon capture and sequestration. Most importantly, by giving a complete answer to the question of existence of solutions, our result enables correct diagnosis of algorithmic failure, problem stiffness and non-convergence in computational algorithms.
@@ 1 Introduction
FLUID flow in pipeline infrastructure networks driven by junction/node potentials through a network consisting of pipelines and pressure regulating devices (like pumps/compressors) is relevant to energy distribution and environmental remediation, such as in delivery of hydrogen/natural gas, water distribution, or transport of carbon dioxide in carbon capture and sequestration. Under steady-state conditions, the governing equations (which are coupled PDEs [1]) reduce to a non-linear system of algebraic equations that relate pressures at junctions/nodes with mass flows in the pipes and other devices throughout the network. The solution of such a non-linear system is a crucial ingredient in various decision support tools for system design and operation that aim to ensure safe and reliable transport of fluids through the network.
@@
Non-linear algebraic systems may have no solution, a unique solution or even multiple solutions, and before designing algorithms to solve such systems, it is important to answer the question of existence and uniqueness of solutions if possible. The question of existence is not an abstruse mathematical curiosity, but one which has fundamental significance from the viewpoint of modelling, control as well as computation. The equations representing the model of a real system are a mathematical construct that aims to mimic the physical response of real systems. While the occurence of a physical phenomena may be undeniable and obvious, the verification that the equations can simulate this phenomena with fidelity is tantamount to proving existence of a solution under certain hypotheses or assumptions. Thus, investigation of existence of a solution is linked to the validity of the model itself. In the control of linear systems, the existence of a solution is rarely spoken about since most linear ODE systems under consideration are almost always regular and guarantee existence and uniqueness. However, in case of nonlinear systems, notions such as controllability cannot be well-defined unless it is known that a solution exists [2]. Finally, from the perspective of computation, existence of a solution is an assurance that if successively finer numerical approximations do not converge, it is the algorithm that needs modification and not the problem.
@@
Notwithstanding the connection to these fundamental issues, rarely do systems arising in practice afford us either the luxury of first proving/determining existence of a solution before turning to algorithms for approximating it, or the courtesy of readily satisfying the hypotheses required for the invocation of some mathematical theorem that can assert the existence of a solution. Thus, research efforts march onwards to develop computational algorithms assuming that a solution exists, with uniqueness of solution lending support when applicable. Such is the case even with a well-established, justly celebrated model like the Navier-Stokes equation for incompressible fluids. The subject of this article is the steady-state equations governing adiabatic, unidirectional fluid flow in networks. The canonical form of these equations corresponds to different contexts, such as natural gas delivery, water distribution, or transport of carbon dioxide in carbon capture and sequestration. However, the case of gas flow is considered here since it is the most complicated of the three due to the compressible nature and equation of state that governs the density.
@@
For gas flow, the non-linear operator has been shown to have a unique solution, if one exists [1, 3, 4, 5], and several attempts have been made to develop efficient computational algorithms [6, 7, 4, 1] agnostic to the grid topology and size that can determine a solution. However, none of these algorithms succeed with large grid sizes, and the absence of a definite result on existence of solutions hobbles further development; for it is not possible to distinguish between algorithm failure and non-existence of a solution.
@@
In this letter, we show that a unique solution exists for such non-linear gas flow systems if the term solution is interpreted in terms of potentials and flows rather than pressures and flows. This result applies to other fluid flow networks as well and is a significant improvement from [1], where the authors conclude that even for the simple case of flow in a single pipe, a solution may not exist. A careful examination reveals that their definition of a solution is in terms of pressures and flows and hence does not contradict the result presented here.
@@
We use topological degree theory and homotopy invariance [8, 9] to assert existence of a solution by starting with a linear system (where the degree computation is trivial) that can be continuously transformed to obtain the non-linear system of interest. There are subtle but crucial differences from homotopic continuation methods [10, 11, 12] which are constructive and postulate a trajectory connecting the solution of a known problem to the solution of the non-linear system of interest. Homotopic continuation is mainly used as a numerical algorithm [13, 14], but theoretical assertions require that the system satisfy very stringent hypotheses that are verifiable only in rare cases [15] where the system has some special structure. In contrast, assertions of existence following degree theory and homotopy invariance as used here, require weaker assumptions that even allow the Jacobian to be singular at a solution.
@@
The article is structured as follows: After setting up the notation and recapitulating the governing equations in Section 2, we state and summarize the propositions and assertions required in Section 3 before using those to prove the main result in Section 4 and concluding with Section 5.
2 Mathematical Preliminaries
@@
Let $G(V,E)$ denote the graph of the pipeline network, where $V$, $E$ are the set of junctions and edges of the network respectively. The edge set is given by $E=P\bigcup C$ where $P$ and $C$ are the set of pipes and compressors respectively. A pipe $(i,j)$ in $P$ connects the junctions $i$ and $j$ with a resistance coefficient $\beta_{ij}>0$. Also, a compressor $(i,j)$ in $C$ connects junctions $i$ and $j$, which we assume are geographically co-located. For each junction $i$, we let $p_{i},q_{i}$ denote the nodal value of the pressure and the injection at $i$ respectively. For any edge $(i,j)$ (pipe or compressor) that connects junctions $i$ and $j$, we let $\phi_{ij}$ denote the steady flow through the edge. Finally, for each compressor $(i,j)$, we use $\alpha_{ij}\geqslant 1$ to denote the specified pressure boost ratio or compressor ratio. It is assumed that the direction of flow is in the direction of pressure boost, i.e., $i\to j$, and hence $\phi_{ij}>0$ for a physically reasonable compressor $(i,j)$. Each junction $i$ is assigned to one of two mutually disjoint sets consisting of the slack junctions $N_{s}$ and non-slack junctions $N_{ns}$ respectively, depending on whether pressure $p_{i}$ or gas injection $q_{i}$ at the node is specified. Thus $V=N_{s}\cup N_{ns}$.
@@
The nature of the gas is prescribed by the equation of state chosen to model the pressure-density relationship of the gas and it determines a potential function $\pi:\mathbb{R}\to\mathbb{R}$ through
[…]
$\pi^{\mathrm{ideal}}(p)=p^{2}/2\;,$ (2a)
$\pi^{\mathrm{cnga}}(p)=b_{1}p^{2}/2+b_{2}p^{3}/3\;,$ (2b)
with
[…]
@@
If for all $x,\alpha\in\mathbb{R}$, $\pi(\alpha x)=g(\alpha)\pi(x)$ for a fixed $g:\mathbb{R}\to\mathbb{R}$, then it is necessary and sufficient that $\pi,g$ have the form $\pi(x)=kx^{r},g(\alpha)=\alpha^{r}$ for some $k,r\in\mathbb{R}$. It is assumed that $\pi,g$ are differentiable and non-constant.
@@
The sufficiency of the stated representation is obvious. For necessity, observe that if $x=0$, $\pi(0)=g(\alpha)\pi(0)$. If $\alpha=0$, $\pi(0)=g(0)\pi(x)$. These two conditions imply that $\pi(0)=g(0)=0$ in order to rule out the trivial case of constant functions. Now consider $\alpha x\neq 0$. Differentiating with respect to $x$ and $\alpha$, one has $\pi^{\prime}(\alpha x)\alpha=g(\alpha)\pi^{\prime}(x)$ and $\pi^{\prime}(\alpha x)x=g^{\prime}(\alpha)\pi(x)$ respectively. Together, they imply $\alpha g^{\prime}(\alpha)/g(\alpha)=x\pi^{\prime}(x)/\pi(x)=r$ (for some $r\in\mathbb{R}$) from which we may conclude that $\pi(x)=kx^{r},g(\alpha)=c\alpha^{r}$ for some $k,c\in\mathbb{R}$. However, this immediately yields $c=1$. ∎
The potential function $\pi=\pi^{\mathrm{ideal}}$ satisfies Lemma 1 (with $k=1/2,r=2$) while $\pi=\pi^{\mathrm{cnga}}$ does not.
The system of equations for the network may now be summarized as follows [1], with the understanding that quantities with a *** superscript constitute given data:
$\mathcal{G}\mathcal{F}_{p}:$ $p_{i}=p^{*}_{i}\;\forall i\in N_{s}.$ (4d)
Definition 1.
If there exists $\bm{p}\in\mathbb{R}^{|V|},\bm{\phi}\in\mathbb{R}^{|E|}$ that satisfies the $\mathcal{G}\mathcal{F}_{p}$ system (4d), the system is solvable, or alternatively, a generalized pressure-solution to system (4d) exists.
Proposition 1.
Given assumptions (A1)–(A3), and a monotonic potential function $\pi$, solutions to the $\mathcal{G}\mathcal{F}_{p}$ system (4d), if they exist, are unique.
Proof.
See [3, 1]. ∎
If we denote a nodal potential $\pi_{i}=\pi(p_{i})\;\forall i\in V$, we can eliminate the pressures in favour of the potentials in (4a) and (4d). This leaves (4b) as the only remaining equation where variables corresponding to nodal pressures are present.
The equation (4b) relating pressures $p_{i},p_{j}$ implies that $\pi(\alpha^{*}_{ij}p_{i})=\pi_{j}$. When the potential function $\pi$ is as in Lemma 1, we obtain an equation relating potentials $\pi_{i},\pi_{j}$ for some suitable $\gamma^{*}_{ij}=g(\alpha^{*}_{ij})$. Thus the system (4d) is transformed into a new system:
@@
Clearly, transforming $\mathcal{G}\mathcal{F}_{p}$ system (4d) to $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) is possible when the ideal gas equation of state is assumed but the CNGA equation of state forbids such a transformation in (4b). Further analysis of the system corresponding to the CNGA equation of state is postponed to Sec. 4.
@@
If there exists $\bm{\pi}\in\mathbb{R}^{|V|},\bm{\phi}\in\mathbb{R}^{|E|}$ that satisfies the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d), the system is solvable, or alternatively, a generalized potential-solution to system (5d) exists.
@@
If the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) is solvable, then the solution is unique.
@@
One can interpret the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) as a special case of $\mathcal{G}\mathcal{F}_{p}$ system (4d) with the equation of state $\rho(p)=1$, and $\pi(p)=p$. By Proposition 1, solutions to the $\mathcal{G}\mathcal{F}_{p}$ system (4d), if they exist, are unique. Then it is clear that if the $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) is solvable, then the solution is unique. ∎
For given Definitions 1 and 2 assuming Lemma 1, note that solvability of $\mathcal{G}\mathcal{F}_{p}$ system (4d) implies solvability of $\mathcal{G}\mathcal{F}_{\pi}$ system (5d). However, the converse is false, for even if system (5d) is solvable, for some $\pi_{i}\in\mathbb{R}$, it is possible that there is no $p_{i}\in\mathbb{R}$ such that $\pi(p_{i})=\pi_{i}$. In any case, the equation (5b) will pose no problems since the form of $\pi$ is dictated by Lemma 1.
Remark 2.
We do not impose any restrictions (such as sign or domain) motivated by physical considerations on $\bm{\pi}$, $\bm{p}$ or $\bm{\phi}$ in Definitions 1 and 2, hence we use the term generalized solution.
@@
  3. (Q3)
    Can we prescribe a domain within which solutions (if they exist) must lie?
We can answer (Q3) for the system (5d) by estimating the upper bound using the given data as follows:
@@
$\beta_{m}\triangleq\max_{(i,j)\in P}\beta^{*}_{ij},\quad\alpha_{m}\triangleq\max_{(i,j)\in C}\alpha^{*}_{ij},$
$\phi_{m}\triangleq\left(\max_{i\in N_{ns}}|q^{*}_{i}|\right)\cdot|N_{ns}|,$
$\pi_{m}\triangleq\left(\max_{i\in N_{s}}|\pi^{*}_{i}|+|P|\beta_{m}\phi_{m}^{2}\right)\cdot(\alpha_{m})^{|C|},$
$L\triangleq\max(\phi_{m},\pi_{m}).$
@@
$\mathcal{D}_{\mathcal{H}}=\{\bm{x}\in\mathbb{R}^{|V|+|E|}:|x_{i}|<L\}$ (6)
@@
The upper bounds above have been derived by considering the maximum possible flow into any single pipe, then calculating the maximum possible pressure by considering a node connected to the slack node with maximum pressure, and furthermore, taking the maximal possible compression achievable in the full network.
In Section 4, we shall find that the answer to (Q1) is that for an ideal gas, $\mathcal{G}\mathcal{F}_{\pi}$ system (5d) has a unique generalized potential solution. If the solution admits non-negative potentials, then the $\mathcal{G}\mathcal{F}_{p}$ system (4d) is also solvable, having at least one generalized pressure solution.
3 Background
Before proving the existence results outlined in the previous section, we state (without proof) some mathematical results that we shall soon invoke.
Definition 3.
A point $\bm{y}$ is a regular value of a map $\bm{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ if $\det(D\bm{f}(\bm{x}))\neq 0\;\forall\;\bm{x}\in\bm{f}^{-1}\{\bm{y}\}$.
Definition 4.
Given a map $\bm{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and an open bounded domain $\mathcal{D}\subset\mathbb{R}^{n}$ such that
  1. 1.
    $\bm{f}$ is continuously differentiable on $\mathcal{D}$,
  2. 2.
    $\bm{y}$ is a regular value of $\bm{f}$,
  3. 3.
    $\bm{y}\notin\bm{f}(\partial\mathcal{D})$,
the degree [8, Chapter 6] of $\bm{f}$ at $\bm{y}$ given $\mathcal{D}$ is defined as
$\mathrm{deg}(\bm{f},\mathcal{D},\bm{y})=\sum_{i}\mathrm{sgn}(\det(D\bm{f}(\bm{x}_{i})))\;\forall\;\bm{x}_{i}\in\bm{f}^{-1}\{\bm{y}\}.$
If $\bm{f}^{-1}\{\bm{y}\}$ is empty we define $\mathrm{deg}(\bm{f},\mathcal{D},\bm{y})=0$.
Remark 3.
The definition of degree can be extended to allow non-regular values without affecting subsequent results, namely the following two propositions. The proofs and other details of degree theory may be found in [8, Chapter 6] or [9].

### 08 Apr 2024 v4

@@ Abstract
The steady-state solution of fluid flow in pipeline infrastructure networks driven by junction/node potentials is a crucial ingredient in various decision-support tools for system design and operation. While the nonlinear system is known to have a unique solution (when one exists), the absence of a definite result on the existence of solutions hobbles the development of computational algorithms, for it is not possible to distinguish between algorithm failure and non-existence of a solution. In this letter, we show that for any fluid whose equation of state is a scaled monomial, a unique solution exists for such nonlinear systems if the term solution is interpreted in terms of potentials and flows rather than pressures and flows. However, for gases following the CNGA equation of state, while the question of existence remains open, we construct an alternative system that always has a unique solution and show that the solution to this system is a good approximant of the true solution. The existence result for flow of natural gas in networks also applies to other fluid flow networks such as water distribution networks or networks that transport carbon dioxide in carbon capture and sequestration. Most importantly, our result enables correct diagnosis of algorithmic failure, problem stiffness, and non-convergence in computational algorithms.
@@ 1 Introduction
FLUID flow in pipeline infrastructure networks driven by junction/node potentials through a network consisting of pipelines and pressure-regulating devices (like pumps/compressors) is relevant in the context of energy distribution and environmental remediation, such as in the delivery of hydrogen/natural gas, water distribution, or transport of carbon dioxide in carbon capture and sequestration. Under steady-state conditions, the governing equations (which are coupled PDEs [1]) reduce to a nonlinear system of algebraic equations that relate pressures at junctions/nodes with mass flows of the pipes and other devices throughout the network. The solution of such a nonlinear system is a crucial ingredient in various decision support tools for system design and operation to ensure the safe and reliable transport of fluids through the network.
@@
Nonlinear algebraic systems may have no solution, a unique solution, or even multiple solutions, and before designing algorithms to solve such systems, it is important to answer the question of existence and uniqueness of solutions if possible. The question of existence is not an abstruse mathematical curiosity but one that has fundamental significance from the viewpoint of modeling, control, and computation. The equations representing the model of a real system are a mathematical construct that aims to mimic the physical response of real systems. While the occurrence of a physical phenomenon may be obvious and undeniable, the verification that the equations can simulate this phenomenon with fidelity is tantamount to proving the existence of a solution under certain hypotheses or assumptions. Thus, investigation of existence of a solution is linked to the validity of the model itself. In the control of linear systems, the existence of a solution is rarely spoken about since most linear ODE systems under consideration are almost always regular and guarantee existence and uniqueness. However, in the case of nonlinear systems, notions such as controllability cannot be well-defined unless it is known that a solution exists [2]. Finally, from the perspective of computation, the existence of a solution is an assurance that if successively finer numerical approximations do not converge, it is the algorithm that needs modification and not the problem.
@@
Notwithstanding the connection to these fundamental issues, rarely do systems arising in practice afford us either the luxury of first proving/determining existence of a solution before turning to algorithms for approximating it, or the courtesy of readily satisfying the hypotheses required for the invocation of some mathematical theorem that can assert the existence of a solution. Thus, research efforts march onwards to develop computational algorithms assuming that a solution exists, with uniqueness of solution lending support when applicable. Such is the case even with a well-established, justly celebrated model like the Navier-Stokes equation for incompressible fluids. The subject of this article is the steady-state equations governing adiabatic, unidirectional111the only non-zero component of flow/velocity is along the length of the edge fluid flow in networks. The canonical form of these equations corresponds to different contexts, such as natural gas delivery, water distribution, or transport of carbon dioxide in carbon capture and sequestration. However, the system of equations corresponding to natural gas subsume the other cases and thus the analysis and conclusions obtained herein for natural gas apply without caveat to the other fluid flow systems.
@@
For (natural) gas flow, the nonlinear system of equations has been shown to have a unique solution, if one exists [1, 3, 4, 5], and several attempts have been made to develop efficient computational algorithms [6, 7, 4, 1] agnostic to the grid topology that can determine a solution. However, none of these algorithms succeed with large grid sizes, and the absence of a definite result on existence of solutions hobbles further development; for it is not possible to distinguish between algorithm failure and non-existence of a solution. Moreover, non-existence of solutions (pressures and flows) has been shown [1] even for the simplest case of flow in a single pipe.
@@
In this letter, we show that for any fluid whose equation of state is a scaled monomial, such as the ideal gas, a unique solution exists for these nonlinear gas flow systems if the term solution is interpreted in terms of potentials and flows rather than pressures and flows as in [1] and this result applies to other fluid flow networks as well.
@@
We use topological degree theory and homotopy invariance [8, 9] to assert existence of a solution by starting with a linear system (where the degree computation is trivial) that can be continuously transformed to obtain the nonlinear system of interest. There are subtle but crucial differences from homotopic continuation methods [10, 11] which are constructive and postulate a trajectory connecting the solution of a known problem to the solution of the nonlinear system of interest. Homotopic continuation is mainly used as a numerical algorithm [12] and requires that the system satisfy very stringent hypotheses that are verifiable only in rare cases [13] where the system has some special structure. In contrast, assertions of existence following degree theory and homotopy invariance as used here require weaker assumptions that even allow the Jacobian to be singular at a solution.
@@
The article is structured as follows: We state and summarize the mathematical propositions and assertions required in Section 2. After setting up the notation and recapitulating the governing equations in Section 3, we prove the main result in Section 4 and conclude with Section 5.
2 Mathematical Background
The following results from topological degree theory and homotopy invariance will be invoked subsequently. Proofs and other details of degree theory may be found in [8, Chapter 6] or [9].
Definition 1.
A point $\bm{y}$ is a regular value of a map $\bm{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ if $\det(\bm{f}^{\prime}(\bm{x}))\neq 0\;\forall\;\bm{x}\in\bm{f}^{-1}\{\bm{y}\}$.
Definition 2.
Given a map $\bm{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and an open bounded domain $\mathcal{D}\subset\mathbb{R}^{n}$ such that (1) $\bm{y}$ is a regular value of $\bm{f}$, (2) $\bm{f}$ is continuously differentiable on $\mathcal{D}$, and (3) $\bm{y}\notin\bm{f}(\partial\mathcal{D})$, the degree of $\bm{f}$ at $\bm{y}$ over $\mathcal{D}$ is defined as
$\mathrm{deg}(\bm{f},\mathcal{D},\bm{y})=\begin{cases}0&\mathrm{if}\;\bm{f}^{-1}\{\bm{y}\}\;\textrm{is empty}.\\ \sum\limits_{n}\mathrm{sgn}(\det(\bm{f}^{\prime}(\bm{x}_{n})))&\forall\;\bm{x}_{n}\in\bm{f}^{-1}\{\bm{y}\}.\end{cases}$
Remark 1.
Proposition 1.
If $\mathrm{deg}(\bm{f},\mathcal{D},\bm{0})\neq 0$, then $\bm{f}$ has a zero/root in $\mathcal{D}$.
Proposition 2 (Homotopy invariance).
Note that the definition of degree can be extended to allow non-regular values in Propositions 1 and 2.
3 Model Description
The adiabatic, unidirectional flow of compressible gas in a single pipeline is described by the simplified Euler equations [14].
$\frac{\partial\rho}{\partial t}+\frac{\partial\phi}{\partial x}=0,\;\frac{\partial\phi}{\partial t}+\frac{\partial p}{\partial x}=-\lambda\frac{\phi|\phi|}{\rho(p)},$ (1)
where $\rho,p,\phi$ are the density, pressure, and mass flux, respectively. The additional parameter $\lambda$ incorporates the friction factor and diameter of the pipe. For steady-state, the system simplifies to
$\phi^{\prime}(x)=0,\;\rho(p)p^{\prime}(x)=-\lambda\cdot\phi|\phi|.$ (2)
The nature of the gas is prescribed by the equation of state $\rho(p)$ chosen to model the pressure-density relationship of the gas and it determines a potential function $\pi:\mathbb{R}\to\mathbb{R}$ through
[…]
which allows (2) to be integrated to the algebraic form
$\phi_{1}=\phi_{2}=\phi_{12},\;\pi(p_{1})-\pi(p_{2})=\lambda\cdot\phi_{12}|\phi_{12}|\cdot(x_{2}-x_{1}).$
$\rho(p)=p,\;\pi^{\mathrm{ideal}}(p)=p^{2}/2\;,$ (4a)
$\rho(p)=b_{1}p+b_{2}p^{2},\;\pi^{\mathrm{cnga}}(p)=b_{1}p^{2}/2+b_{2}p^{3}/3\;,$ (4b)
[…]
In most cases, the equation of state $\rho(p)$ is a polynomial, and so the potential function also has a polynomial form.
@@ 3 Model Description
In order to describe the flow equations for a network of pipelines, we first introduce some notation. Let $G(V,E)$ denote the graph of the pipeline network, where $V$ and $E$ are the set of junctions and edges of the network, respectively. The edge set is given by $E=P\bigcup C$ where $P$ and $C$ are the set of pipes and compressors, respectively. A pipe $(i,j)$ in $P$ connects the junctions $i$ and $j$ with a resistance coefficient $\beta_{ij}>0$. Also, a compressor $(i,j)$ in $C$ connects junctions $i$ and $j$, which we assume are geographically co-located. For each junction $i$, we let $p_{i},q_{i}$ denote the nodal value of the pressure and the injection at $i$, respectively. For any edge $(i,j)$ (pipe or compressor) that connects junctions $i$ and $j$, we let $\phi_{ij}$ denote the steady flow through the edge. Finally, for each compressor $(i,j)$, we use $\alpha_{ij}\geqslant 1$ to denote the specified pressure boost ratio or compressor ratio. It is assumed that the direction of flow is in the direction of pressure boost, i.e., $i\to j$, and hence $\phi_{ij}>0$ for a physically reasonable compressor $(i,j)$. Each junction $i$ is assigned to one of two mutually disjoint sets consisting of the slack junctions $N_{s}$ and non-slack junctions $N_{ns}$ respectively, depending on whether pressure $p_{i}$ or gas injection $q_{i}$ at the node is specified. Thus $V=N_{s}\cup N_{ns}$.
The system of equations for the network may now be summarized as follows [1] (*** superscript indicates given data):
$\mathcal{G}\mathcal{F}_{p}:$ $p_{i}=p^{*}_{i}\;\forall i\in N_{s}.$ (5d)
Definition 3.
Any $\bm{p}\in\mathbb{R}^{|V|},\bm{\phi}\in\mathbb{R}^{|E|}$ that satisfies the $\mathcal{G}\mathcal{F}_{p}$ system (5d) is called a generalized pressure-solution.
The definition of generalized pressure-solution allows the pressures and flows to take any real values. However, due to physical considerations we make a distinction between the equations being solvable and feasible. As outlined in [1], for both the ideal and the non-ideal gas (with CNGA equation of state), a feasible solution will have non-negative pressures and non-negative flow in compressors. Otherwise, the solution/system is infeasible.
@@
Proposition 3.
Given assumptions (A1)–(A3), and a monotonic potential function $\pi$, a generalized-pressure solution to the $\mathcal{G}\mathcal{F}_{p}$ system (5d), if it exists, must be unique.
Proof.
See [3, 1]. ∎
@@
If $\pi,g$ are real-valued, differentiable functions and for all $x,\alpha\in\mathbb{R}$, $\pi(\alpha x)=g(\alpha)\pi(x)$, then it is necessary and sufficient that $\pi$ be a scaled monomial, i.e., $\pi(x)=kx^{r}$ for some $k,r\in\mathbb{R}$.
@@
The sufficiency of the stated representation is obvious. For necessity, observe that if $x=0$, $\pi(0)=g(\alpha)\pi(0)$. If $\alpha=0$, $\pi(0)=g(0)\pi(x)$. These two conditions imply that either $\pi,g$ are constants or $\pi(0)=g(0)=0$. Now consider $\alpha x\neq 0$ for which $\pi(x)\neq 0,g(\alpha)\neq 0$. Differentiating with respect to $x$ and $\alpha$, one has $\pi^{\prime}(\alpha x)\alpha=g(\alpha)\pi^{\prime}(x)$ and $\pi^{\prime}(\alpha x)x=g^{\prime}(\alpha)\pi(x)$ respectively. Together, they imply $\alpha g^{\prime}(\alpha)/g(\alpha)=x\pi^{\prime}(x)/\pi(x)=r$ (for some $r\in\mathbb{R}$) from which we may conclude that $\pi(x)=kx^{r}$ for some $k\in\mathbb{R}$. ∎
If we denote a nodal potential $\pi_{i}=\pi(p_{i})\;\forall i\in V$, we can eliminate the pressures in favour of the potentials in (5a) and (5d). The equation (5b) relating pressures $p_{i},p_{j}$ implies that $\pi(\alpha^{*}_{ij}p_{i})=\pi_{j}$.
Remark 2.
When the potential function $\pi$ is a scaled monomial as in Lemma 1, we obtain an equation relating potentials $\pi_{i},\pi_{j}$ for some suitable $\gamma^{*}_{ij}=g(\alpha^{*}_{ij})$ and the system (5d) is transformed into a new system:
@@
The potential function $\pi=\pi^{\mathrm{ideal}}$ satisfies Lemma 1 (with $k=1/2,r=2$) while $\pi=\pi^{\mathrm{cnga}}$ does not. Thus $\mathcal{G}\mathcal{F}_{p}$ system (5d) transforms to $\mathcal{G}\mathcal{F}_{\pi}$ system (6d) for the ideal gas equation of state but not for the CNGA equation of state. This will be investigated further in Sec. 4.
@@
A $\bm{\pi}\in\mathbb{R}^{|V|},\bm{\phi}\in\mathbb{R}^{|E|}$ that satisfies the $\mathcal{G}\mathcal{F}_{\pi}$ system (6d) is called a generalized potential-solution.
@@
Given assumptions (A1)–(A3), a generalized-potential solution to $\mathcal{G}\mathcal{F}_{\pi}$ system (6d), if it exists, must be unique.
@@
One can interpret the $\mathcal{G}\mathcal{F}_{\pi}$ system (6d) as a special case of $\mathcal{G}\mathcal{F}_{p}$ system (5d) with the equation of state $\rho(p)=1$, and $\pi(p)=p$. By Proposition 3, solutions to the $\mathcal{G}\mathcal{F}_{p}$ system (5d), if they exist, are unique. ∎
For given Definitions 3 and 4 assuming Lemma 1, note that solvability of $\mathcal{G}\mathcal{F}_{p}$ system (5d) implies solvability of $\mathcal{G}\mathcal{F}_{\pi}$ system (6d). However, the converse is false, because for some $\pi_{i}\in\mathbb{R}$, there may be no $p_{i}\in\mathbb{R}$ such that $\pi(p_{i})=\pi_{i}$.
@@
For now, we can only prescribe a domain within which solutions for the system (6d) (if they exist) must lie. This we do by estimating upper bounds using the given data as follows:
@@
$\beta_{m}\triangleq\max_{(i,j)\in P}\beta^{*}_{ij},\;\alpha_{m}\triangleq\max_{(i,j)\in C}\alpha^{*}_{ij},\;\phi_{m}\triangleq\max_{i\in N_{ns}}|q^{*}_{i}|\cdot|N_{ns}|,$
$\pi_{m}\triangleq\left(\max_{i\in N_{s}}|\pi^{*}_{i}|+|P|\beta_{m}\phi_{m}^{2}\right)\cdot(\alpha_{m})^{|C|}.$
@@
$\mathcal{D}_{\mathcal{H}}=\{\bm{x}\in\mathbb{R}^{|V|+|E|}:|x_{i}|<\max(\phi_{m},\pi_{m})\}$ (7)
@@
These bounds have been derived by considering the maximum possible flow into any single pipe, then calculating the maximum possible pressure by considering a node connected to the slack node with maximum slack pressure, and furthermore, taking the maximal possible compression achievable in the full network.
In Section 4, we shall find that the answer to (Q1) is that the system for an ideal gas has a unique generalized potential solution and if the solution admits non-negative potentials, then it has a generalized pressure solution as well.
4 Results and Implications
The stated facts in Section 2 will now be put to use. Before that however, we record the following result (restated with slight modification from [1]) that guarantees the invertibility of a linear system with particular structure and is essential for the subsequent analysis.